OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -8 and a6 = -8,192? -327,680 -393,216 -458,752 -524,288

OpenStudy (anonymous):

can you please provide me with the correct answer and how to solve it. thanks.

hartnn (hartnn):

an= a1.r^{n-1} so a6=(-8)r^5 find r from here.

OpenStudy (anonymous):

a1= a = -8 a6= a.r^6-1 = -8192 -8.r^-5 = -8192 r= \[\sqrt[5]{8192/8} = \sqrt[5]{1024}= 4\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

9th term will be a.r^8 u have a and r values...

OpenStudy (anonymous):

so would the answer be -393,216